# HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models

### v2

@@ 4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training portion of the test datasets, including the Implicit_Hate set. To simulate low-resource settings, we choose 256 samples uniformly at random from each class, then augment them with explanations as described in Section 3.1. We perform K-shot sample (K $\in\{32,64,128,256\}$) to finetune both Base and Pretrained versions of the models.
@@
Interestingly, decoder-only models (Llama) considerably outperform encoder-decoder COT-T5-XL on the 2 and 3-way classification tasks, yet the reverse is observed for the nuanced 6-way Implicit_Hate. In fact, only COT-T5-XL consistently scales with the increment in training data to reach max F-1 score of 0.56, while Llama models plateau even at K=256 on this task at sub-0.3 range.
We further select the best performing model at K=256 for each dataset and finetune their Base versions with the entire training data, without explanations and definitions. As shown in Figure 2(b) and Table 4, in-domain finetuning with HateCOT enables models to attain competitive results even with only a fraction of the training data. Furthermore, prompting for explanations enables Llama 13B and COT-T5-XL to attain performance surpassing using the full training data for Latent_Hate and Implicit_Hate.
@@ 4.3 In-context Learning
As an alternative to in-domain fine-tuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template as in Figure 5. We run inference for classification results over 5 random seeds, which also randomly permutes their order.
Figure 2(c) shows the mean, minimum and maximum values macro F-1 scores over the seeds for the base, HateCOT-tuned only (pre-trained), and in-domain finetuned (K=256) versions of Llama 7B and Llama 13B. COT-T5-XL regularly generates overly repetitive outputs, and thus omitted. The range of F-1 scores are large regardless of settings, an observation in line with the variance of in-context learning in literature Lu et al. (2022); Dong et al. (2022). Unsurprisingly, base models’ performances tend to be inferior to their tuned counterpart. Finetuned models display superior performance than the other 2 counterparts on average. Interestingly, the max F-1 scores of finetuned models in ICL are not appreciably better than those in the zero-shot counterparts (Figure 2(b)). In contrast, except for Llama 7B on HateXplain, the best scores of pretrained models approach those of the finetuned models–particularly for Llama 13B.
This finding suggests another advantage of pre-training on HateCOT then maximize performance via ICL without additional in-domain finetuning, an area that has attracted growing attention Min et al. (2021); Wang et al. (2023); Ye et al. (2023b). Nevertheless, there exists the trade-off: ICL examples with explanations extend significantly the context length, and ICL inferencing takes considerably more time compared to zero-shot, making the latter more resource-efficient overall.
@@ 4.4 Assessment and Recommendations
    Practitioners should choose model of sufficient number of parameters for the task. Larger instruction-tuned LLMs appear to more effectively capitalize on HateCOT pretraining regimen before in-domain finetuning.
@@
GPT-4 Human (Frequency Count) Human (Majority Vote)
Model Base Pretrained $p$-val. Base Pretrained Base Pretrained
COT-T5-XL 27 (5.4%) 473 (94.6%) *** 62 (24.8%) 188 (75.2%) 6 (12%) 44 (88%)
Llama 7B 109 (21.8%) 391 (78.2%) *** 109 (43.6%) 141 (56%) 19 (38%) 31 (62%)
Llama 13B 122 (24.4%) 378 (75.6%) *** 114 (45.6%) 136 (54.4%) 22 (44%) 28 (56%)
Table 3: Comparison of Base and Pretrained models in GPT-4 and Human Evaluation. *** indicates $p$-value $\ll$ 0.05 by the Wilcoxon rank test. Output preferred by at least 3 out of 5 annotator wins the Majority Vote for that sample.
@@ 5 Quality of Explanations
We analyze the quality of the explanations generated by various versions of 3 models COT-T5-XL, Llama 7B, Llama 13B through evaluation by both human annotators and OpenAI’s GPT-4 following similar literature Liu et al. (2023); Li et al. (2023). An additional assessment of Target Identification is presented in Appendix A.5.
@@
We randomly select 500 samples per model whose explanations from both the Base and HateCOT-pretrained versions agree on the predicted label from all combined test sets. Using the template in Figure 6, we ask GPT-4 to select its preferred explanation for the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. We present the explanations from these 2 versions in a randomized order for the same post. Using a similar setup, we recruit 13 annotators from the crowdsource platform Amazon Mechanical Turk and solicit their rating on 50 samples per model (Appendix A.4). Each post is annotated by 5 humans, resulting in 750 annotations across 3 models.
Converting GPT-4’s ratings to an ordinal scale (0 if Base model is preferred, 1 otherwise), we perform the Wilcoxon rank sum test with the null hypothesis $H_{0}$ (no significant difference between the number of times Pretrained or Base model is preferred over the other). In Table 3, we observe the preference for explanations of the Pretrained models across all settings, even when considering the human tally by using majority vote ($\geq$ 3 out of 5 annotations). The human preference margin is smaller for Llama 13B Pretrained, likely due to this model’s already strong generative capabilities.
@@ 5.2 QA 2: Inter-model Comparison
Inspired by Wang et al. (2023); Lin et al. (2023); Yang et al. (2023), we assess the quality of explanations generated by finetuned models (K=256) on the following criteria:
  • $\triangleright$
    Informativeness: whether the model provides factually correct and relevant information in its explanation.
@@
Using the template in Figure 7, we collect numerical ratings from GPT-4 on a scale from 1 (least) to 5 (very) on these criteria for 200 posts per model per dataset, for a total of 2,400 samples. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets for each model. Ratings for Informativeness appear consistently lower than those of the other criteria, likely due to the challenging context-dependent nature of generating post-specific relevant information. Overall, Persuasiveness and Soudness ratings around 4 out of 5 suggest high likelihood of acceptance by human users for content moderation. We also note the general correlation between the models’ better classification performance (Figure 2(b)) and higher mean ratings on each dataset, a useful artifact to calibrate models.
@@ 6 Conclusion
We show empirically that our HateCOT dataset considerably enhances offensive speech detection even with limited training data while producing high-quality explanations. We invite researchers to build upon our work to address other related low-resource areas.
7 Limitations
@@ A.1 Data Pre-Processing
Dataset Best Model @ K =256 F1 Retrained F1 Base + Full Retrained F1 % Retrained Data Size Full Training Data Size Retrained Data %
HateCheck LLAMA 13B 0.95 0.99 96% 512 1,864 27%
HateXplain LLAMA 13B 0.64 0.72 89% 768 12,088 6%
Latent_Hate LLAMA 13B 0.66 0.64 103% 768 11,460 7%
Implicit_Hate COT-T5-XL 0.56 0.38 147% 1,536 2,707 57%
Table 4: Comparison of performance metrics for the best performing model at K=256 in-domain finetuning vs. finetuned on the full training set, without explanations or definitions. Retrained F1% denotes the percentage of macro F-1 score on the K=256 retrained model over that of the version trained on full data. Similarly, Retrained Data % denotes the percentage of data size used by the Retrained model over the size of the full data.
@@ A.3 Variants of Prompt Template for Explanation
A.4 Human Annotation for QA Experiment
@@
The general demographic breakdown of the 13 annotators are described below:
@@
A.5 QA 3: Target Identification
Finally, we investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have the Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 8. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
@@ A.5 QA 3: Target Identification
Sap et al.,2019 Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.

### v4

@@ 4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training portions of the test datasets, including Implicit_Hate. To simulate low-resource settings, we choose 256 samples uniformly at random from each class, then augment them with explanations as described in Section 3.1. Both the Base and Pretrained versions of the models are then finetuned using various K-shot $\in\{32,64,128,256\}$ training data from this pool.
@@
Interestingly, decoder-only models (Llama) considerably outperform encoder-decoder COT-T5-XL on the 2 and 3-way classification tasks, yet the reverse is observed for the nuanced 6-way Implicit_Hate. On this task, only COT-T5-XL consistently scales with the increment in training data to reach the max F1 score of 0.56, while Llama models plateau at sub-0.3 range even at K=256.
We further select the best performing model at K=256 for each dataset and finetune their Base versions with the entire training data and no explanations no definition. In Figure 2(b) and Table 3, in-domain finetuning after HateCOT achieves competitive results even with only a fraction of the full training data. Furthermore, prompting for explanations enables Llama 13B and COT-T5-XL to attain performance that surpasses using the full training data on Latent_Hate and Implicit_Hate.
@@ 4.3 In-context Learning
As an alternative to in-domain finetuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template shown in Figure 5. We run inference for classification results over 5 seeds, which also randomly permutes their order.
Figure 2(c) shows the mean, minimum and maximum values of macro F1 scores over the seeds for the base, HateCOT-pretrained only (Pretrained), and in-domain finetuned (K=256) versions of Llama 7B and Llama 13B. COT-T5-XL regularly generates overly repetitive outputs, and thus omitted. The range of F1 scores is large regardless of settings, an observation in line with the variance of in-context learning in literature Lu et al. (2022); Dong et al. (2022). Unsurprisingly, base models’ performances tend to be inferior to their finetuned counterpart. Interestingly, the max F1 scores of finetuned models with ICL are not appreciably better than those in the zero-shot counterparts (Figure 2(b)). In contrast, except for Llama 7B on HateXplain, the best scores of pretrained models approach those of the finetuned models–particularly for Llama 13B.
This finding suggests another advantage of pre-training on HateCOT: boosting performance via ICL without additional in-domain finetuning, an area that has attracted growing attention Min et al. (2021); Wang et al. (2023b); Ye et al. (2023b). Nevertheless, there exists the trade-off: ICL examples with explanations extend significantly the context length, and ICL inferencing takes considerably more time compared to zero-shot, making the latter more resource-efficient overall.
Dataset Best Model @K=256 F1 @K=256 F1 Base + Full F1 % Data Size @K=256 Data Size Full Data %
HateCheck LLAMA 13B 0.95 0.99 96% 512 1,864 27%
HateXplain LLAMA 13B 0.64 0.72 89% 768 12,088 6%
Latent_Hate LLAMA 13B 0.66 0.64 103% 768 11,460 7%
Implicit_Hate COT-T5-XL 0.56 0.38 147% 1,536 2,707 57%
Table 3: Comparison of performance metrics for the best performing models finetuned using K=256 in-domain post HateCOT vs. finetuned on the full training set and no explanation nor definition. F1% denotes the percentage of macro F1 score of the K=256 finetuned model over that of the model trained on full data. Similarly, Data % denotes the percentage of data size used by the K=256 regimen over the full data.
@@ 4.4 Assessment and Recommendations
    The most consistent benefit of HateCOT is its capacity to enable data-efficient in-domain finetuning following pretraining.
  • $\triangleright$
    Practitioners should choose models of sufficient number of parameters for the task. Larger instruction-tuned LLMs appear to more effectively capitalize on HateCOT pretraining regimen before in-domain finetuning.
@@ 5 Quality of Explanations
In addition to the enhanced classification prowess, we investigate whether pretraining on HateCOT also improves the quality of explanation LLMs. To this end, the following 2 human quality assurance (QA) experiments are conducted. In QA 1, we assess if human annotators prefer the explanations generated by the base or HateCOT-pretrained LLMs (COT-T5-XL, Llama 7B, Llama 13B). In QA 2, we perform in-domain K-shot finetuning on the aforementioned models and examine how annotators evaluate their generated explanations. An additional assessment of Target Identification is presented in Appendix A.6.
@@
From the outputs of the test sets generated by the 3 LLMs, we select 50 samples uniformly at random whose explanations from the Base and HateCOT-pretrained versions agree on the predicted label, for a total of 150 samples and 300 explanations. We then recruit 13 annotators from the crowdsource platform Amazon Mechanical Turk and solicit their annotation on these explanations (Appendix A.5). Using the template in Figure 6, we ask the annotators to indicate their preferred explanation that better suits the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. Each post is annotated by 5 humans, resulting in 750 annotations.
In Table 4, we observe that the raw frequency count for explanations generated by the HateCOT-pretrained models exceed their base version’s. Similarly, even when tallying by majority vote–where the explanation is chosen by at least 3 out of 5 annotators–preference for those generated by the Pretrained models still prevails. We note that the preference margin is smaller for Llama 13B Pretrained, likely due to this model’s already strong generative capabilities.
Human (Frequency Count) Human (Majority Vote)
Model Base Pretrained Base Pretrained
COT-T5-XL 62 (24.8%) 188 (75.2%) 6 (12%) 44 (88%)
Llama 7B 109 (43.6%) 141 (56.4%) 19 (38%) 31 (62%)
Llama 13B 114 (45.6%) 136 (54.4%) 22 (44%) 28 (56%)
Table 4: Comparison of Base and Pretrained models in Human Evaluation. Frequency Count : count per annotation; Majority Vote indicates aggregate count by the version is preferred by at least 3 out of 5 annotators.
@@ 5.2 QA 2: Inter-model Comparison
Inspired by Wang et al. (2023a); Lin et al. (2023); Yang et al. (2023), we assess the quality of explanations generated by finetuned models (K=256) on the following criteria:
@@
For this QA task, we recruit 6 annotators also from Amazon Mechanical Turk (Appendix A.5). Using the template in Figure 7, we collect their numerical ratings on a scale from 1 (least) to 5 (very) on these criteria for 50 posts per model per dataset, for a total of 600 annotations. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets. The average Ovearall ratings for both Persuasiveness and Soudness are above 3.2 out of 5, indicating generally positive reception by human evaluators. Interestingly, there exists a degree of correlation between the models’ better classification performance (Figure 2(b)) and higher mean ratings on each dataset, a useful artifact to calibrate models. These ratings may serve as benchmarks for future works.
@@ 6 Conclusion
We show empirically that our HateCOT dataset considerably enhances offensive speech detection even with limited training data while producing high-quality explanations. We invite future research to explore other benefits of LLM-augmented data and extend them to other related low-resource areas.
Limitations
@@
Acknowledgement
This work is funded by the University of Maryland’s Institute for Trustworthy AI in Law & Society (TRAILS). We thank the service of ACL ARR reviewers, area chairs and the editors of the EMNLP conference for our paper’s publication.
@@ A.3 Variants of Prompt Template for Explanation
A.4 Quality Review of Sample Explanations
Elaborating on the criteria outlined in Section 3.1, we review the quality of the generated explanation by GPT-3.5-Turbo on the following items:
  • Grammatically correct
  • Succinct in their justification of the chosen label.
  • Persuasive and logical in their reasoning for the chosen label
We discard explanations that are too verbose, and/or not choosing the label already provided by human annotation (fortunately, this scenario happens rarely, likely due to the presence of existing rationales guiding the extra generated outputs). We also remove explanations that conjure incorrect and/or irrelevant facts to the context to discourage hallucination and encourage a high degree of pertinent to the target post.
@@ A.4 Quality Review of Sample Explanations
A.5 Human Annotation for QA Experiment
A.5.1 QA 1
@@
The demographic breakdown of the 13 annotators are described below:
@@
A.5.2 QA 2
For this experiment, each annotation task consists of the explanations generated by 3 models are grouped by the sample. This division resutls in 200 tasks. 6 Amazon Mechanical Turk workers are recruited, with similar qualification criteria as described above.
The demographic breakdown of the 6 annotators are described below:
  • Gender: Female (3), Male (3)
  • Age: 18-29 (1), 30-39 (2), 40-49 (3)
  • Education: 2-year college (2), 4-year college (4)
A.6 QA 3: Target Identification
We investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have the Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 8. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
@@ A.6 QA 3: Target Identification
Sap et al.,2020 Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.